bill533:

23 meters is the same as how many feet? Hint: 1 m ≈ 3.3 ft Round your answer to the nearest tenth. ________

SmokeyBrown:

Let's see, 23 * 3.3 is 75.9, so that'd be the number of feet in 23 meters

bill533:

75.9 correct ?

SmokeyBrown:

Yup. Since there are 3.3 feet in 1 meter, there would be 3.3*23 feet in 23 meters. And that's equal to 75.9

bill533:

Find the complex conjugate of the given number. -10 Write your answer in the form a+b𝑖.

SmokeyBrown:

Oh, haha, I don't really know what that means, sorry

bill533:

can you look it up

SmokeyBrown:

Ok, so a complex conjugate of a complex number has the same value in the real number but the opposite value of the imaginary number. So, the complex conjugate of something like 3 + 4i would be 3 - 4i The real number, 3 stays the same, and the imaginary number, 4i, is multiplied by -1

SmokeyBrown:

Now, in the question, we're given -10, but that's just a real number, with no imaginary part. So, if I'm understanding the concept correctly, the complex conjugate of -10 would be -10

bill533:

A 2-cup bottle of hand soap costs $4.80. What is the price per fluid ounce? $_____

SmokeyBrown:

So, there's 8 fluid ounces in a cup. That means 2 cups is 16 fluid ounces. So, if the soap costs $4.80 for 16 fluid ounces, 1 ounce would cost 4.80/16, which is $0.30

bill533:

Every day, Katie walks by the same grocery store on her way home from school. Most days she goes in to buy a snack. Today she's craving peanut-butter pretzels, so she fills a small bag from the bulk bin section. After weighing the bag on the scale, she removes 0.15 pounds of pretzels. Now the bag weighs 0.25 pounds. Which equation can you use to find the weight w of the bag before Katie removes the pretzels? A. w – 0.25 = 0.15 B. w – 0.15 = 0.25 C. w + 0.15 = 0.25 D. 0.15w = 0.25 Solve this equation for w to find the weight of the bag before Katie removes the pretzels. ______ pounds

SmokeyBrown:

So, if Katie removes 0.15 from the original weight, w, and gets 0.25 as a result, the equation we're looking for would be B

bill533:

Solve this equation for w to find the weight of the bag before Katie removes the pretzels. ______ pounds

SmokeyBrown:

You'd want to solve for w by getting w by itself on one side of the equation

SmokeyBrown:

Your final result would be w=___

bill533:

w = 25

bill533:

am i right ?

SmokeyBrown:

No, sorry, I'm not sure how you got that. Starting from the equation, w-0.15=0.25 We want to get w on its own, so we'd add 0.15 to both sides in order to cancel the -0.15. That gives us w -0.15 +0.15 = 0.25 +0.15 Which simplifies to w = 0.40

bill533:

https://www.desmos.com/calculator/ooem3vasws https://www.desmos.com/calculator/ooem3vasws btw do these two look alike to you ?

SmokeyBrown:

Yeah, I think they are similar. If you multiply the lengths of the smaller triangle by... 3, I think, it'll be equal to the large triangle

SmokeyBrown:

Yeah, that looks good to me

bill533:

ok thank you

bill533:

1 attachment